dy/dx + 4y= 8



1
Expert's answer
2022-06-03T02:52:20-0400
"dy\/dx=-4y+8"

"\dfrac{dy}{y-2}=-4dx"

Integrate


"\int\dfrac{dy}{y-2}=-\int4dx"

"\ln|y-2|=-4x+\ln C"

"|y-2|=Ce^{-4x}"